You are here: Foswiki>Commissioning Web>Logs>09May2012_Log (08 May 2012, JuanGuerra)Edit Attach
.---+++ UT 20120508 AO#2 Commissioning

Observer: SEsposito, RBriguglio, APuglisi, FQuiros-Pacheco (LBTO)
Telescope Operator: DGonzalez (LBTO)
Additional AO Support: JCGuerrar
Telescope Support: JUrban(LBTO)
Instrument Support: TShih (LBTO)

Summary

Installation of the RR to work in a new bin recosntructor. Reaquire pupil and reconstructor for Bin1.

Test of the reconstructor during the night. We have seen the intermittent failure of the use of this reconstructor. The secondary RIP with the inner ring jumped out. Seeing variable 1.1"-1.7". Secopndary telemetry dayta are saved for post-processing analysis and investigation of the problem.

Details

Daytime activity to work in a new reconstructor for bin1. This new reconstructor will be tested during the night.

New reconstructor for bin1

seeing variable 1.1"-1.5"

We can see that the new and old reconstructor are giving some problems. The inner ring of actuator jump up and force the shell to RIP. Many presets have been sent and secondary telemetry data are saved. analysis of the data didn't reveal a clear issue which trigger those secondary events. An issue is the pupil separations that change during night time and daytime RR. Daytime RR is 36 pixel separation adn during the night is 35.5 or 35.7 pixels.

At the end of the night we went up to the W-Unit to check is something is loose the board, but seems everything is tide and solid.

finish 10:30

-- JuanGuerra - 07 May 2012
Topic revision: r2 - 08 May 2012, JuanGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback